Suppose the market demand curve is given by Qd = 80 - 10P, and the market supply curve is given by Qs = 10 + 15P. What is the equilibrium price and quantity?
Answer #1

The equilibrium occurs at the point where the quantity demanded is equal to the quantity supplied. So, there is neither shortage or surplus quantity in the market.

In order to calculate the equilibrium price and equilibrium quantity, the demand and supply must be equated.

Equating demand and supply:

80 - 10P = 10 + 15P

25P = 70

P = 70 / 25

P* = 2.80

Putting P* = 2.80 in the supply function:

Q = 10 + 15 (2.80)

Q = 10 + 42

Q* = 52

So, the equilibrium quantity is 52 units and the equilibrium price is 2.80.